OpenStudy (anonymous):

When I was considering what to do with my $10,000 Lottery winnings, my broker suggested I invest half of it in gold, the value of which was growing by 9% per year, and the other half in CDs, which were yielding 5% per year, compounded every 6 months. Assuming that these rates are sustained, how much will my investment be worth in 10 years? (Round your answer to the nearest cent.)

OpenStudy (anonymous):

gold: FVg = 5000(1+0.09)^10 CD: FVc = 5000(1+0.05/2)^20 Calculate each and add together.

OpenStudy (anonymous):

so its 25103.29

OpenStudy (anonymous):

no sorry i forgot to divide

OpenStudy (anonymous):

20029.89

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

= 11836.82 + 8193.08 = 20029.90